OpenStudy (anonymous):

Please help will fan and medal!!! Choose the point-slope form of the equation below that represents the line that passes through the points (-3,2) and (2,1) A) y+3=-5(x-2) B) y-2=-5(x+3) C) y+3=-1/5(x-2) D) y-2=-1/5(x+3)

OpenStudy (anonymous):

@Hotchellerae21 do you think you could help me with this one too? Its different than the last since it doesn't give you the slope?

OpenStudy (anonymous):

Would I still use y=mx+b?

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1)? The equation is of the form y - y1 = m(x - x1) (x1, y1) is either of the given points; "m" is the slope. Find the slope. m = (y2-y1)/(x2-x1) ....= (1-2)/(2+3) = -1/5 However, this slope is not in any of the options, and I checked my work, so either you mistyped something, maybe one of the points, or the people making the book made a mistake. Plug in either point for y1 and x1 in the equation. Apparently they used (-3, 2). y - 2 = -(1/5)(x - (-3)) y - 2 = -(1/5)(x + 3) Given the equation y − 3 = 1/2 (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form? Slope-intercept form is "y=mx+b", "m" being the slope and "b" being the y-intercept. Simplify the given equation to this form. y - 3 = (1/2)x + 3 y = (1/2)x + 6 Again, none of the options given are the correct answer. Given the equation y − 3 = 1/2 (x + 6) in point-slope form, identify the equation of the same line in standard form? Standard form is of the form "Ax + By = C". y - 3 = (1/2)x + 3 y - (1/2)x = 6 Eliminate the fraction by multiplying both sides of the equation by 2. 2y - x = 12 Switch the x and y terms (the order) and multiply both sides so that the first term is positive. x - 2y = -12 (answer A) Give medal please if that is what your looking for thanks

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

Ive mistyped something?

OpenStudy (anonymous):

@shinalcantara could you please help me... Im not sure how to do this

OpenStudy (anonymous):

@ganeshie8 Im stuck do you think you could help me please? Im really not sure what to do and was hoping someone could explain it to me

OpenStudy (anonymous):

Oh Nevermind I get it @openkkkkk thank you so much :D

OpenStudy (anonymous):

welcome can u give medal

OpenStudy (anonymous):

Already did :3

OpenStudy (anonymous):

thanks!!!!